OMC Transmission measurement after the 2nd deep cleaning
The 2nd deep cleaning didn't improve the transmission. (See Attachment 2)
The measured loss was 0.044+/-0.002
The four cavity mirrors in OMC #1 were each scrubbed using acetone and a cotton swab.
Then, the four mirrors were painted with First Contact (picture attached). The First Contact was allowed to dry for 20 minutes, then removed while using the top gun.
Conclusion on the cleaning of OMC #001
- After a couple of first contact cleaning trials and deep cleaning, the total loss was measured to be 0.045+/-0.004.
This indicated a slight improvement from the loss measured at LLO before any cleaning (0.064+/-0.004).
However, the number did not improve to the level we marked in 2013 (0.028+/-0.004).
- This loss level of 4.5% is comparable to the loss level of OMC #3, which is currently used at LHO.
Therefore, this OMC #1 is still a useful spare for the site use.
- Some notes / to-do regarding this unit:
1) The beam dump with melted black glass was removed. A new beam dump needs to be bonded on the base.
2) The connector bracket still needs to be replaced with the PEEK version.
3) The PZT of CM1 has been defunct since 2013. Combining LV and HV drivers is necessary upon use at the site. (LLO used to do it).
- Removed the first contact we left on Monday.
- Measured transmission (Set1) Very high loss! Total optical loss of 18.5%! Observation with the IR viewer indicated that CM1 has bright scattering. We suspencted a remnant of FC.
- Applied the second FC on the four cavity mirrors. This made the CM1 sport darker.
- Measured the transmission (Set1~Set3). We had consistent loss of 4.2~5.0%. We concluded that this is the limitation of this OMC even with the cleaning.
Photo of the BS1 AR cleaning process
Attachment 1: Before cleaning. Foggy surface is visible.
Attachment 2: After FC cleaning. The structure of the deposited material is still quite visible.
Attachment 3: Acetone scrubbing. Cotton Q-tip was used so that the stick does not melt with acetone.
Attachment 4: After acetone scrubbing. Nicely clean!
Acetone scrubbing was applied to HR/AR of BS1, FM1, FM2, BS2, and HR of CM1 and CM2. (total 10 surfaces)
Then final FC paint was applied to these 10 surfaces.
We'll come back to the setup on Thu for FC peeling and loss measurement.
The damaged black glass was removed from the OMC breadboard leaving the glass base.
The black glass pieces were bonded very tightly on the FS base with EP30-2. The apparent amount of the bond was not so much but it was such hard that removal by hand was not possible.
We decided to give drips of Acetone on the base hoping the gradual dissolving of EP30-2. Using a knife edge, the "filets" of the bonds were removed, but the BD was still tight.
By wedging the black glass-black glass bonding with the nife edge, the left side (the directly damaged one) was taken off from the structure leaving a tiny fragment of the glass on the base.
The remaining one was even stronger. We patiently kept dripping Acetone on the base and finally, the black glass piece was knocked off and removed from the base.
Attachment 1: The base right after the black glass removal.
Attachment 2: The black glass pieces were stored in a container with Al foil + clean cloth bed. The damaged and fogged surfaces faced up.
Attachment 3: The zoom-in shot of the black glass pieces.
Attachment 4: The base was wiped with Acetone and cleaned with FC. We will bond another BD assembly on the base, presumably using the UV epoxy.
Monday, November 14, 2022
Camille and Koji did a "deep cleaning" of OMC#1:
1) Applied First Contact to the mirror surfaces. Removed first contact after ~10 minutes.
2) Acetone scrub of the mirror surfaces with a cotton swab.
3) Applied First Contact again. Removed after ~10 minutes. We left the FC paint on for the work on Thu.
The foggy spot on the input mirror was unchanged after the first round of First Contact. But the foggy spot came off during the acetone scrub.
The measured total optical loss of the OMC was
1st: 0.015 +/- 0.003
2nd: 0.085 +/- 0.005
3rd: 0.0585+/- 0.0008
4th: 0.047 +/- 0.002
In avegrage the estimated loss is
Loss = 0.055 +/- 0.014
This is unchanged from the measurement at LLO after the FC cleaning
Loss = 0.053 +/- 0.010
Friday, Nov 11th, 2022
Setting up OMC #1 for transmission measurements:
The laser beam was aligned to the OMC cavity. The OMC cavity was locked and the transmission measurements were recorded.
Clean Supply Ordered
Fiber matching: 43.2/56.7 = 76%
S/P-pol ratio 0.7/43.2 = 1.6%
Began setting up fiber assembly for OMC testing:
-Aligned fiber mount to maximize transmission through fiber
-Adjusted polarization at output of fiber to minimize s-polarized output.
fiber input: 56.7 mW
fiber output:43.2 mW
s-polarized output: 700 uW
Upon the LLO work, the current PD arrangement in the cages are:
B1 OMC1 PDT (A1-23)
B2 OMC1 PDR (A1-25)
B3 original (C1-03)
B4 OMC2 PDT (B1-22)
C1 OMC2 PDR (B1-23)
C2 original (C1-08)
C3 original (C1-09)
C4 original (C1-10)
Started July 15, 2022 and finished Aug 30. So it took ~1.5 months (with a couple weeks of break)
Class B special tools
First Contact Kit
Bonding kit (excl EP30-2 bond)
Power meters (excl Power meter controller)
Cable bracket replacement kit
Optics / Optomechanics
=== Action done on Aug 30 ===
Fiber MM setup / Fiber coupler mount
Glass Beamdumps (for optical testing)
Thorlabs fiber coupler tool
General bent nose plier for fiber
Thorlabs collimator tiny allen
Spare High QE PDs
Spare OMC bags / Zip bags
Balance Mass 10g Qty 8 (Different Type D11*** 1.25" dia), 20g Qty 10 / Mass damper D1700301 -04 / Mass damper screws SHCS 1/4-20 x 1.25 Qty 25 / 1" screws and 1 1/8" screws
Shipping request: https://services1.ligo-la.caltech.edu/FRS/show_bug.cgi?id=25002
=== Low supply! ===
Here is the balance mass info for the LLO OMC#001 analyzed from the photographs
If we attach the additional mass, longer 1/4-20 screws (1", 1" 1/8, 1" 1/4) are going to be used.
Inspected the past LLO add-on mass configuration.
There are unknown masses at the DCPD side. It looks like a small SS mass with an estimated mass of 5g. But the DCC number is unknown.
We are going to add 10g on each corner as well as the damping aterial. We should be able to figure out the fastener / mass configuration.
The OMC #002 is ready for shipment.
Attachment 1: Work done on Sept 19, 2022
Other attachments: Putting the OMC in the pelican case.
EP30-2 bonding setup
Inspection of the delaminations in the optics side
Inspection of the bonding on the suspension interface side. All look good.
#1 The Invar bar on the cable bracket (DCPD side)
Added short (frosted) Al bars (Attachment 1) to the short sides of the invar bar. (Attachments 2/3). Some glue was sucked into the delamination gap by capillary action (=good) (Attachment 4)
#2 The Invar bar on the cable bracket (QPD side)
Added short (frosted) Al bars to the short sides of the invar bar. (Attachments 3/5). Maybe some glue was sucked into the delamination gap??? Not so clear. (Attachment 4)
#3 The Invar bar reinforced in 2016
Added a short (frosted) Al bars to a short side of the invar bar (Attachment 6). On both sides of the 2016 reinforcement, rectangular prisms are added (Attachment 6)
Some capillary action is visible beneath the invar bar (Attachment 7)
Leave it as it is for a day
Checked the delamination status:
New cable ties were installed on the cable pegs attached to the long sides of the cable bracket.
The cable bracket was successfully replaced.
- The hole size extension is going forwared now.
- Madeline and Chub are cleaning (sonicating) a drill (29/64=0.4531")
- The parts in a bag were brought to the 40m C&B lab.
- The hole is going to be 11mil=0.28mm larger than the recommendation (0.442").
It's not a D-hole. The connector has a rounded-rectangular flange that fits into the PEEK parts.
So I don't think it's an issue.
- Chub has a proper spanner to fasten the nuts. We want to use it here and LLO.
Now we got the C&Bed parts to continue to work on the cable bracket replacement.
1) Helicoil insertion
1/4-20 Helicoils were inserted into the 6 thread holes of D1300052. It went mostly okay. We witnessed that the Helicoil insertion tool delaminated the plating of the Helicoils upon insertion (Attachment 1). Stephen mentioned that this is not usual, but we didn't find anything further such as increased friction, more debris, etc. So we decided to go forward.
2) EP30-2 Kit
The EP30-2 kit was transferred from the 40m clean room to the OMC lab. The EP30-2 kit tracking was updated via C1900343
3) D1300052 reinstallation -> FAIL
Now resumed to the installation of D1300052 bracket. However, the hole size of the bracket is just a bit too small compared with the size of the mighty mouse connectors. It was already quite tight with the metal version. However, this PEEK version seems to have 0.1 mm further small diameter, and then the connectors do not penetrate the holes. The plan could be
1) Use a razor blade to shave the hole inner circle.
2) Use a cleaned drill bit to make the hole size 0.2mm bigger.
Checking the spare parts
- Conclusion for OMC#2: need PEEK cable ties
- for more OMCs: need more BHCS / PEEK cable ties / Helicoils
- (Attachment 1) The connector nut rings were removed using an angled needle nose plier. The connector shell has a tight dimension relative to the hole on the bracket. But of course, they could be extracted.
- The 4 screws mounting the bracket to the invar blocks were successfully removed. No extra damage to the bonding.
- (Attachment 2) The plan was to remove the cable pegs by unfastening the button head 1/4-20 screws from the bracket and then just replace the bracket with the new one. However, these screws were really tight. The two were successfully removed without cutting the PEEK cable ties. Two cable ties were necessary to be cut to detach the bracket+pegs from the fragile OMC. Then one screw was removed. However, the final one could not be unfastened. This is not a problem as we are not going to recycle the metal cable bracket... as long as we have spare parts for the new bracket.
- (Attachment 3) Right now, the new bracket is waiting for the helicoils to be inserted. So the OMC lid was closed with the cables piled up. Just be careful when the lid is open.
- D1300052-V3 SN001 is going to be used (Attachment 1)
- This is the PEEK version of the cable bracket (Attachment 2). The side thread holes have no Helicoils inserted. This needs to be done!
Connector arrangement check / cable routing check
Attachment 3: Connector Arrangement from the Northside
Attachment 4: Connector Arrangement from the South side
Attachment 5: Cable routing (Northside down)
At this point, the delamination of the V shape beam dumps was visible. This is the subject of bonding reinforcement.
The optical surfaces were coated with FirstContact to keep them clean / somewhat protected during the transportation.
The PD aperture was sealed with FirstContact "caps" (made by Kate in 2016?).
- DLPCA-200 trans-impedance amplifier was calibrated.
Keithley source meter 2450 was connected to the amp. Provide current and read the output voltage with the precision digital voltage meter (Agilent/Keysight).
Gain: 999.7V/A@7mA, 999.6V/A@8mA
- From the power meter spec, Thorlabs S401C seemed the best (+/-3%). So the QEs of the 9 PDs were checked with this power meter again.
- All PDs exhibited the QE of 0.95~0.96. It's all relative as the power meter has a systematic error.
- Tried to clean B1-22 and B1-23 PDs. They didn't show significant improvement after the cleaning. To avoid the unnecessary risk of damaging the PDs, further cleaning was not performed. (Some photos were attached)
- What we can do is use this result as the relative measurements.
- For OMC#2, B1-22 is the DCPD(T) and B1-23 is the DCPD(R). C1-03 and C1-12 are the spares, according to this latest result.
- At LLO, we track down the source of the throughput reduction (-10%). The QEs of the PDs are going to be tested in the same setup at once to compare their PDs and our PDs.
- 4 CLASS A wire clamp obtained from the OMC spare
- 4 more DIRTY wire clamp obtained from WB experiments (they no longer use these)
Once the later ones are C&Bed, we have enough.
- Installed the High QE PDs to OMC #002
Upon the installation, the legs of the PDs were cut by 3mm. Also, the tab of the PD could not be embedded in the DCPD housing. Therefore, the tabs were cut.
The alignment looked just fine. The weak reflections are directed to the black glass beam dumps.
- After the installation, the QEs were measured.
It is so confusing. So I decided to make the QE test setup.
Ophir RM9 with chopper (+/-5%): 8.97mW
Thorlabs S140C integrating sphere (+/-7%): 9.11mW
Thorlabs S130C PD power meter (+/-7%): 9.15mW
Thorlabs S401C thermal power meter (+/-3%): 8.90mW
So there looks ~3% discrepancy between S130C and S401C
Then tried to measure the QE of C1-03@Cage B3 with Ophir RM9
- Initial state: QE=0.95
- First FirstContact application: QE went up to 0.973
- Second FirstContact application: QE = 0.974, basically no change
- Calibrate the trans-impedance amp with Keithley
- Apply FC to B1-22 and B1-23 to see if there is an improvement
- The power should be measured with S401C because the accuracy seems better (+/-3%).
- Take photos of the PD FC process
General To Do:
- Backscatter test 2nd trial
- Start applying the first contact to the optical surfaces
- Beam dump cleaning
- Apply FC cap to the PDs
- Delamination repair (light side)
- Delamination repair (dark side)
- Cable bracket replace (dark side)
More epoxy delamination check:
DCPD R (Attachment 1): Found half delaminated
DCPD T (Attachment 2): Found half delaminated
QPD1/QPD2 (Attachment 3): Looks fine
In total we need to fix bonding of three invar bases (including the one for the cable bracket)
OMC Reinforcement blocks
1. P/N D1600316; Version v4; Type 01; Qty 30; Source Chemistry Machine Shop
2. P/N D1600316; Version v4; Type 02; Qty 15; Source Chemistry Machine Shop
3. P/N D1600316; Version v4; Type 01; Qty 40; Source Resource MFG PO S422806
4. P/N D1600316; Version v4; Type 02; Qty 40; Source Resource MFG PO S422806
Stephen asked Srinath for the ICS entry.
Stephen made the C&B request https://cleanandbake.ligo.caltech.edu/clean_and_bake/request/1708/
Madeline was asked to take care of the C&B.
Also, the Torr Seal box was returned to Madeline.
SRS LCR meter SRS720 was returned to Downs as before.
Measure the power ratio between the forward-propagating and reverse-propagating beams.
- To increase the incident laser power, NPRO Current ADJ was set to be 0 (increased from -50)
- 1st: Without the baffle 0.373 +/- 0.001 uW / With the baffle 0.318 +/- 0.001 uW
- 2nd: Without the baffle 0.370 +/- 0.001 uW / With the baffle 0.318 +/- 0.001 uW
- 3rd: Without the baffle 0.370 +/- 0.001 uW / With the baffle 0.317 +/- 0.001 uW
==> 53.3 +/- 0.6 nW
- The main transmission was 84.0mW
==> Backpropagation ratio was 0.634+/-0.007 ppm
- Direct measurement of the OMC was after BS 96.6mW
==> Backpropagation power from the cavity: 61.3 +/- 0.7 nW
- Cavity transmission for the matched beam is Tcav RinputBS = 0.963
==> Incident resonant TEM00 power 100.3mW
- Reflection 61.3+/-0.7 nW x RinputBS = 60.8+/-0.7 nW
-> The effective reflectivity for the mode-matched resonant TEM00 beam incident on the OMC (1st steering mirror) is 0.606+/-0.007 ppm
The profile of the beam incident on the fiber input
The fiber input was deflected by a 45deg mirror. The beam profile was measured with WincamD. The beam was too strong (~60mW) even at the smallest pump power (ADJ -50) of the NPRO. So the two ND20 filters were added to the lens right before the 45 deg mirror and the camera.
The measured profile had some deviation from the nice TEM00 particularly around the waist. This could be a problem of the too small beam on the ND filter and the CCD.
This is not an issue as we just want to know the approximate shape of the beam.
For the fiber coupling, if we have the beam waist radius of ~200um it is sufficient for decent coupling.
The backscatter beam is supposed to appear in the backpropagation path. The transmission of the OMC has a couple of optics, it's not easy to access that beam.
To try to deflect the beam either horizontally or vertically, small optical pieces were made. (Attachment)
These are the combination of the optics
- Thorlabs PF05-03 Fused Silica Mirror Blank (dia12.7mm x t 6.0mm) + Thorlabs 1/2"sq BB Dielectric Mirror BBSQ05-E03
- Thorlabs PF05-03 Fused Silica Mirror Blank (dia12.7mm x t 6.0mm) + Thorlabs ME05-G01 Protected Al Mirror (dia12.7mm x t 3.2mm) + Thorlabs MRA10-K13 Right-Angle Prism Nd:YAG 10mm
Torr seal was used as the bonding epoxy. It uses a 1:2 volume mixture (not easy because of the viscosity) and is relatively fast to cure (in a couple of hours).
The test piece showed some softness after 3~4 hours so I left the parts cured overnight at room temp (i.e. 18degC)
Bond reinforcement blocks for the invar brackets:
The Windows laptop for WincamD/Beam'R2 (DELL Vostro3300) was not functional.
- Windows 7 got stuck in the starting up process (Google "startup repair loop")
- The battery can't charge and the adapter connection is flaky
I decided to newly install Win10.
I made a new bootable Win10 DVD from the ISO downloaded from IMSS. The ISO file was converted to CDR using Disk Utility on Mac.
This deleted the past disk partitions. The installation process has no trouble and Win10 ran successfully. The machine is slow but still acceptable for our purpose.
Dataray Version 7.1H25Bk was downloaded from the vendor website https://dataray.com/blogs/software/downloads and installed successfully.
The devices ran as expected by connecting the heads and selecting the proper device in the software.
Then, the Win10 fell into "Hibernation Loop" and "Shutdown loop" (after disabling hibernation in the safe mode).
This is probably the combination of extremely slow windows update (feature update i.e. beta OS update) and the occasional shutdown due to the flakiness of the AC connection
Win10 was reinstalled and automatic Win update was disabled via windows policy manager or something like that. Still, it tries to download and update some of the updates (what's happening there!?
Here are my strong recommendations on how to use this laptop
o Power Budget after FirstContact cleaning (2022/07/20)
NPRO ADJ -50 (min)
Fiber incident --.-mW
Fiber output --.-mW
Matching to the fiber ??%
DCPD T = 8.62 +/- 0.01 mW
REFPD = 3.549 +/- 0.001 V
DCPD R = 9.46 +/- 0.01 V
REFPD = 3.562 +/- 0.001 V
CM1 = 74.5 +/- 0.1 uW
REFPD = 3.585 +/- 0.001 V
CM2 = 81.7 +/- 0.1 uW
REFPD = 3.585 +/- 0.001 V
vOFS = -6.197 +/- 0.001 mV (beam blocked)
LOCKED = 47.6 +/- 0.2 mV
REFPD = 3.596 +/- 0.003 V
UNLOCK = 2.700 +/- 0.003 V
REFPD = 3.590 +/- 0.001 V
P_Inc = 19.36 +/- 0.001 mW
REFPD = 3.594 +/- 0.001 V
- Cavity coupling 0.980 (2.0% junk&sidebands)
- Cavity R&T: R=229ppm, T=0.970 (previous T=0.946, 2.4% UP!)
- OMC Throughput (Cavity T x First BS R): T=0.963
- Cavity loss per mirror 42.8 ppm / Round Trip Loss 238ppm
o Power Budget (2022/07/18)
NPRO ADJ -50 (min)
Fiber incident 62.8mW
Fiber output 45.1mW
Matching to the fiber 72%
DCPD T = 8.90 +/- 0.01 mW
REFPD = 3.760 +/- 0.001 V
DCPD R = 8.82 +/- 0.01 V
REFPD = 3.760 +/- 0.001 V
CM1 = 81.4 +/- 0.1 uW
REFPD = 3.767 +/- 0.001 V
CM2 = 86.6 +/- 0.1 uW
REFPD = 3.767 +/- 0.001 V
OFS = -6.214 +/- 0.001 mV (beam blocked)
LOCKED = 57.5 +/- 0.5 mV
REFPD = 3.970 +/- 0.003 V
UNLOCK = 2.816 +/- 0.003 V
REFPD = 3.943 +/- 0.001 V
P_Inc = 20.04 +/- 0.01 mW
REFPD = 3.946 +/- 0.001 V
- Cavity coupling 0.989 (1.1% junk&sidebands)
- Cavity R&T: R=756ppm, T=0.946
- OMC Throughput (Cavity T x First BS R): T=0.939
- Cavity loss per mirror 90 ppm / Round Trip Loss 432ppm
== Initial Preparation ==
== Measurements ==
== Repair / Preparation ==
== Shipping ==
- The lab is chilly (18degC)
- Cleaned the lab and the optical table a bit so that the delicate work can be done. The diode test rig (borrowed from Downs - see OMC ELOG 408 and OMC ELOG 409) was removed from the table and brought to the office (to return on Monday)
- The rack electronics were energized.
- The OMC mirrors in use were returned to the cases and stored in the plastic box.
- The optical table was also cleaned. Removed the old Al foils. The table was wiped with IPA
- The OMC #4 was moved to the other part of the table, and then OMC #2 was placed in the nominal place (Attachment 1). Note that the "legs" were migrated from #4 to #2. There are three poles that defines the location of the OMC Transportation
- The lid was removed and the OMC was inspected (Attachment 2). Immediately found some more delamination of the epoxy beneath the cable bracket (Attachment 3). This needs to be taken care of before shipment.
- The cavity was already flashing as usual, and a bit of alignment made the TEM00 flashing.
- The locking was a little tricky because the LB unit seemed to have a gain-dependent offset. After some adjustment, robust locks were achieved. The cavity was then finely adjusted. Attachment 4 shows the CCD image of the reflection. The core of the spot is more or less axisymmetric as usual. There is also a large helo around the spot. I was not aware of this before. I may need to wipe some of the mirrors of the input path.
- As the satisfactory lock was achieved, I called a day by taking a picture of the table (Attachment 5).
The table width was an inch too large compared to the door width. We need to tilt the table and it seemed too much for us. Let's ask the transportation for handling.
Photo courtesy by Juan